OpenStudy (anonymous):

could someone please add this system of equations by addition method: x-5y=-4 -6x-5y=-46

OpenStudy (karatechopper):

Ah, Elimination. Pick, do you want the x's or the y's to be eliminated?

OpenStudy (anonymous):

the x's! @karatechopper

OpenStudy (karatechopper):

Ok! So, in order to get rid of the x's lets see your two victims. x and -6x. What do you think you should do to make all of the x's go away COMPLETELY?

OpenStudy (anonymous):

hmm, i do not have any idea, I suck at math :/ @karatechopper

OpenStudy (karatechopper):

x-5y=-4 -6x-5y=-46 6x-30y=-24 -6x-5y=-46 -35y=-70 y=2 Now use substitution y=2 x-5y=-4 you can really chose any equation. x-5(2)=-4 5 x 2=10 x-10=-4 add 10 on both sides. x-10=-4 +10 +10 ----------- x=6 Therefore x=6, y=2

OpenStudy (karatechopper):

Understood?

OpenStudy (karatechopper):

Ask any questions you have please :)x.